Fromme's Frocks has the following machine hours and production costs for the last six months of last year:

Month Machine
Hours
Production
Cost
July 16,500 $ 12,150
August 15,000 12,300
September 13,000 9,650
October 17,000 12,230
November 16,300 11,767
December 13,600 9,967

If Fromme expects to incur 12,100 machine hours in January, what will be the estimated total production cost using the high-low method?
Answer #1

Variable cost = (Cost at highest activity - Cost at lowest activity) /(Highest activity-Lowest activity)

= (12,230-9,650)/(17,000-13000) = 0.645

Fixed cost = 12,230 - (17000*0.645) = 1265

Total Cost for 12100 mh

= (12100*0.645)+1265

= 9070 (select the nearest answer)